OpenStudy (anonymous):

Find the area of the shaded region.The graph to the right depicts IQ scores of Adults (85-120 is the shaded region), and those scores are normally distributed with a mean of 100 and a standard deviation of 15

OpenStudy (perl):

you want to find the area between x = 85 to x = 120

OpenStudy (perl):

first change to z score z1 = ( 85 - 100 ) / 15= -1 z2 = ( 120 - 100 ) / 15= 1.33333 P ( -1 < Z < 1.3333) = you can use a z table or calculator